Graph f(x) = –x – 2.
Please help me with this. I dont understand these at all..

Answer by Fombitz(32388)

f(x) = –x – 2.
A function is just a formula where you put a number in and get a number out. Let's generate some values.
For x=0, f(0)= -0-2=-2
For x=1, f(1)= -1-2=-3
For x=2, f(1)= -2-2=-4
Let's do some negative values too.
For x=-1, f(1)= -(-1)-2=-1
For x=-2, f(1)= -(-2)-2=0
Now we have pairs of points (x,y) that describe this function. A grid define the x axis as the horizontal axis and y axis as vertical axis. Plot points on the grid to make your graph.
Let point A be (0,-2)

Then draw a line to connect all the points. Actually your function would continue in a straight line to the edges of your graph paper.